OpenStudy (anonymous):

Simplify (5x+10)/(5x+3) a. (x+7)/x b. 10/3 c. (x+10)/(x+3) d. (5x+10)/(5x+3)

OpenStudy (aum):

Distribute.

OpenStudy (anonymous):

(5x+10)/(5x+3)=((5x+3)+7)/(5x+3)=1+7/(5x+3)